# Akaki Mamageishvili Offchain Labs Benny Sudakov ETH Zurich

### v3

@@ Abstract
We compare the total capital efficiency of secure restaking and Proof-of-Stake (PoS) protocols. First, we consider the sufficient condition for the restaking graph to be secure. The condition implies that it is always possible to transform such a restaking graph into separate secure PoS protocols. Next, we derive two main results: upper and lower bounds on the required extra stakes to add to the validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow as a square root of the number of validators. We also study a complementary question of aggregating secure PoS protocols into a secure restaking graph and provide lower and upper bounds on the PoS savings.
@@ 1 Introduction
    We study a complementary question of aggregating given secure PoS protocols into a secure restaking graph. When aggregating, each validator adds stakes across projects that they secure in separate PoS protocols, resulting in a restaking graph. Similarly to the restaking savings, we define PoS savings as the total additional stake required to add validators in the aggregate restaking graph to make it secure. We first construct an example showing that the resulting restaking graph is not secure, in fact, at least an additional stake equal to the original total stake is required to make it secure, thus showing a lower bound of $1$ on PoS savings. We also provide an upper bound on PoS savings as a function of the underlying restaking graph, which matches a lower bound example for a special set of graphs.
@@ A PoS protocol centered at $s\in S$ is secure implies that
We obtain the following lower bound on this value, by constructing a family of graphs $G$ with an increasing number of services for which $PoSS(G)$ converges to $1$.
@@
Consider PoS protocols with $m$ services. Each service $s\in S$ has value $\pi_{s}=1$ and a security parameter $\alpha_{s}=\frac{1}{m+1}$. There are two types of validators, labeled $s$ for each service $s\in S$ and the last validator labeled m+1{m+1}. Stakes are equal to $\sigma_{s}^{s}=1$ and $\sigma_{m+1}^{s}=\frac{1}{m}$. Note that all $m$ PoS protocols are secure:
  • A validator labeled $s$ does not attack, since it is losing a stake equal to $1$, the value it obtains from attacking a service.
  • A validator labeled m+1{m+1} can not attack the service since its stake share is equal to $\frac{\frac{1}{m}}{1+\frac{1}{m}}=\frac{1}{m+1}$. This is the same as a security parameter of the service.
We claim that the minimum additional stake required to make aggregate restaking graph secure is m1m-1. In the resulting restaking protocol, the validator m+1{m+1} has stake equal to $m\frac{1}{m}=1$. It can attack all services unless its own stake becomes $m$ or one of the other validators gets an extra m1m-1 stake, in which case the claim is trivially true. The extra amount m1m-1 is enough. Consider adding stake m1m-1 to the validator m+1{m+1} so that it now controls the stake amount $m$. This validator does not attack services since it is losing at least the amount it can gain by attacking. Other validators are not able to attack their corresponding services, and it is also not profitable to attack.
Hence, $PoSS\geq\frac{m-1}{m+1}$ for any $m$, since the original total stake amount among the validators is $m+1$. Taking $m\rightarrow\infty$ gives $PoSS\geq 1$. ∎
@@
Note that the upper bound in Proposition 6 matches the lower bound limit obtained in Proposition 5, for graphs where the maximun number of validators over services is equal to two.
@@ 4 Conclusion
In the complementary direction, we demonstrate that aggregating secure PoS protocols into a secure restaking graph can incur a significant overhead. We provide a constant lower bound of $1$ on PoS savings (the extra stake needed post-aggregation to restore security) and non-constant upper bound equal to maximum service degree in the restaking graph.
@@
Two important open questions remain. First is closing the gap between the constant lower bound and the non-constant upper bounds on PoS savings. Second is designing a generic, polynomial-time-checkable sufficient condition for restaking graph security that admits graphs with strictly positive restaking savings.

### v4

@@ Abstract
We compare the total capital efficiency of secure restaking and Proof-of-Stake (PoS) protocols. First, we consider the sufficient condition for the restaking graph to be secure. The condition implies that it is always possible to transform such a restaking graph into separate secure PoS protocols. Next, we derive two main results: upper and lower bounds on the required extra stakes to add to the validators of the secure restaking graph to be able to transform it into secure PoS protocols. In particular, we show that the restaking savings compared to PoS protocols can be very large and can asymptotically grow as a square root of the number of validators. We also study a complementary question of aggregating secure PoS protocols into a secure restaking graph and provide matching lower and upper bounds on the PoS savings.
@@ 1 Introduction
    We study a complementary question of aggregating given secure PoS protocols into a secure restaking graph. When aggregating, each validator adds stakes across projects that they secure in separate PoS protocols, resulting in a restaking graph. Similarly to the restaking savings, we define PoS savings as the total additional stake required to add validators in the aggregate restaking graph to make it secure. We first construct an example showing that the resulting restaking graph is not secure. In fact, at least an additional stake equal to the original total stake times the number of projects/validators minus one is required to make it secure, thus showing a lower bound of number of projects/validators minus one on PoS savings. We also provide an upper bound on PoS savings, which exactly matches a lower bound.
@@ A PoS protocol centered at $s\in S$ is secure implies that
We obtain the following lower bound on the $PoSS$.
@@
Suppose there are $m$ services, each with a value of $\pi_{s}=1$. Suppose there are $m$ validators, and assume each validator secures all $m$ services. For each service $s$, there is its corresponding dedicated validator $s$ with a stake of $\sigma_{s}^{s}=1$, and the other m1m-1 validators a stake of $\epsilon$ each. Set the security parameter for each service to $\alpha_{s}=\frac{(m-1)\epsilon}{1+(m-1)\epsilon}$. First, we check the security of these PoS protocols. An attacking subset containing the dedicated validator would lose at least 1, which results in no profit. A subset excluding the dedicated validator has a maximum combined stake of $(m-1)\epsilon$, which gives a stake fraction of exactly $\alpha_{s}$. Because an attack requires a stake fraction strictly greater than $\alpha_{s}$, this subset cannot attack. Therefore, the independent PoS protocols are secure. The initial total stake is $T(\sigma)=m(1+(m-1)\epsilon)$.
When aggregated into a restaking graph, each validator’s total stake becomes $1+(m-1)\epsilon$. Because all validators secure all services, the total stake on any given service is $m(1+(m-1)\epsilon)$. Consequently, the fraction of the total stake controlled by a single validator on any service is exactly $\frac{1}{m}$. If we choose $\epsilon$ to be sufficiently small (specifically $\epsilon<\frac{1}{(m-1)^{2}}$), then $\frac{1}{m}$ becomes strictly greater than $\alpha_{s}$. This means any single validator has enough stake fraction to independently attack any service. Since this applies to all services, every validator can profitably attack all $m$ services, gaining a value of $m$ while only losing a stake of roughly 1.
To secure this aggregated graph, we must eliminate this profitable attack for all $m$ individual validators. For any given validator, we can remove the attack in one of two ways:
  • Increase the total stake on enough services to force the validator’s stake fraction to drop below $\alpha_{s}$.
  • Increase the validator’s total stake to $\geq m$, so that attacking m services is no longer profitable.
Because $\epsilon$ can be chosen to be arbitrarily close to 0, blocking the attack via the $\alpha_{s}$ threshold requires an arbitrarily large amount of extra stake (inversely proportional to $\epsilon$) to inflate the denominator of the attacking fraction. Therefore, the strictly cheaper option is to increase every validator’s stake to $m$. The extra stake required for each validator is $m-(1+(m-1)\epsilon)$. Across all $m$ validators, the total extra stake $E$ is $m(m-1-(m-1)\epsilon)$.
Then, PoSS for this instance is the ratio of the extra stake to the initial total stake:
${PoSS}(G)=\frac{E}{T(\sigma)}\geq\frac{m(m-1-(m-1)\epsilon)}{m(1+(m-1)\epsilon)}.$
Taking the limit as $\epsilon$ approaches 0, this ratio simplifies exactly to m1m-1. ∎
The lower bound can also be interpreted as n1n-1, since by construction the number of validators is equal to the number of services. This stands in contrast to the upper bound obtained for restaking savings, which was of order $\Theta(\sqrt{n})$. Similarly to Corollary 2, we obtain:
Corollary 3.
$PoSS=\infty.$
@@
Note that the upper bound in Proposition 6 exactly matches the lower bound obtained in Proposition 5, unlike the case of restaking savings, in which case the lower and upper bounds are only asymptotically matching.
@@ 4 Conclusion
In the complementary direction, we demonstrate that aggregating secure PoS protocols into a secure restaking graph can incur a significant overhead. We provide matching lower and upper bounds on PoS savings (the extra stake needed post-aggregation to restore security), equal to maximum service degree in the restaking graph minus one.
@@
One important open question remains. It is designing a generic, polynomial-time-checkable sufficient condition for restaking graph security that admits graphs with strictly positive restaking savings.